OpenStudy (anonymous):

(3x+1)1/3 = -1 (Solving Radical equations) Find the real solution of each equation.

OpenStudy (anonymous):

(3x + 1) = -3 3x = -4 x = -4/3

OpenStudy (anonymous):

Where did the 1/3 go?

OpenStudy (anonymous):

multiply both sides by 3

OpenStudy (anonymous):

Multiply by 3 both Sides

OpenStudy (anonymous):

1/3 Vanishes!

OpenStudy (anonymous):

If I multiply by 3 to both sides how will I show that in work form?

OpenStudy (anonymous):

3(3x +1)1/3= 3(-1)